OpenStudy (anonymous):

Divide a^3 + c^3 by a + c.

OpenStudy (anonymous):

See use the formula: \[a^3 + c^3 = (a+c)(a^2 + c^2 - ac)\] Put this in above expression and tell what you will get??

OpenStudy (anonymous):

So... One?

OpenStudy (anonymous):

\[\frac{(a + c) \times (a^2 + c^2 - ac)}{a +c} = ?\]

OpenStudy (anonymous):

Okay. a^3 / a + c

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

Well, wouldn't it be a x a^2 + a x c^2 + a x -ac and then you simplify and are left with just a^3 and the a + c it was over?

OpenStudy (anonymous):

See (a+c) and (a + c) gets cancel then what are you left with??

OpenStudy (anonymous):

a^2 + c^2 - ac ...Derp.